how do I build up fractions like 5/7 = ?/98 do not understand how to do this please explain
Found 3 solutions by stanbon, richwmiller, solver91311:
Answer by stanbon(75887)

how do I build up fractions like 5/7 = ?/98 do not understand how to do this please explain
-------------------------
5/7 = ?/98
-----------
What was 7 multiplied by to become 98?
Answer: 7x = 98
x = 14
--------------------
So, multiply numerator and denominator of 5/7 by 14 to get:
(5/7)(14/14) = 70/98
------------------------------
Cheers,
Stan H.

Answer by richwmiller(17219)

5/7=x/98
7x=5*98
x=5*98/7
x=70
5*14/7*14=70/98
5/7=70/98


Answer by solver91311(24713)



There are a couple of ways to look at this.

First is you can consider having multiplied by 1 in the form of , like this:



to obtain some number over 98. That means that must equal 98. And that means you can determine the value of by dividing 98 by 7. Once you have the value of you can multiply that by 5 to get the numerator you don't know.

The other way is to look a this like a proportion:



Cross-multiply:



and then just solve for which is your missing numerator.



John